OpenStudy (anonymous):

Find the perimeter of the equilateral triangle JKL with JK = 12

OpenStudy (anonymous):

equilateral triangle has 3 equal sides,a. then , perimeter of the equilateral triangle=3a=3*12=36

OpenStudy (anonymous):

jk=12 p=3s =3(12) =36

OpenStudy (anonymous):

thank you!(: